# Exploring Semantic Labeling Strategies for Third-Party Cybersecurity Risk Assessment Questionnaires

### v1

@@ Abstract
Exploring Semantic Labeling Strategies for Third-Party Cybersecurity Risk Assessment Questionnaires
Ali Nour Eldin  and Mohamed Sellami  and Walid Gaaloul SAMOVAR, Télécom SudParis, Institut Polytechnique de Paris, Palaiseau, France {name}.{last_name}@telecom-sudparis.eu
@@ 1 Introduction
As organizations increasingly rely on third-party providers such as cloud platforms and SaaS vendors, Third-Party Risk Assessment (TPRA) has become a core component of cybersecurity risk management and regulatory compliance. TPRA typically relies on structured questionnaires derived from standards such as ISO/IEC 27001 and NIST 8; National Institute of Standards and Technology (2020). Organizations select relevant questions from internal repositories based on a supplier’s scope and risk profile, evaluate the supplier’s responses, and decide whether to proceed with onboarding or request remediation Keskin et al. (2021); Osunji (2021).
While significant effort is devoted to analyzing supplier responses, a major operational bottleneck arises earlier in the process: selecting relevant questions. Large organizations maintain extensive repositories of compliance questions accumulated across standards and assessment campaigns, yet these repositories are typically unstructured and curated manually Naji et al. (2025). As a result, question selection is repetitive, time-consuming, and difficult to scale.
Most automated approaches frame question selection as a retrieval problem and rely on text similarity Wang and Dong (2020), including cosine similarity over dense embeddings Steck et al. (2024). However, such methods operate purely on semantic similarity and lack explicit knowledge of the underlying control domain (e.g., access control, incident response) or assessment scope (e.g., existence verification, enforcement on critical systems). When these aspects are implicit, retrieval often yields overly generic questions that do not align with the assessment scope.
An alternative is to enrich compliance repositories with semantic labels capturing both control domain and scope, enabling label-based retrieval. However, most repositories lack such labels, and there is no established methodology for deriving them automatically. Existing work has explored automated labeling using topic modeling Jelodar et al. (2019) and clustering-based methods Grootendorst (2022). More recently, Large Language Models (LLMs) have shown strong performance as annotators and can generate human-readable labels without predefined taxonomies Zhang et al. (2023); Tan et al. (2024). Nevertheless, using LLMs as the primary labeling mechanism for large and evolving repositories is costly and sensitive to prompt variability.
@@ 2.1 Third-Party Risk Assessment Systems
TPRA is commonly implemented within third-party risk management (TPRM) workflows that standardize vendor onboarding, questionnaire-based assessments, remediation tracking, and continuous monitoring Keskin et al. (2021). Commercial platforms support these workflows through curated question libraries, predefined mappings to compliance frameworks, and reporting pipelines.
@@ 2.2 Similarity-Based Methods
Similarity-based retrieval ranks questions or controls by textual relevance. Traditional lexical methods depend on vocabulary overlap and are sensitive to phrasing variation Robertson and Zaragoza (2009). Neural retrieval methods embed queries and documents into dense vector spaces, enabling semantic similarity matching Reimers and Gurevych (2019); Karpukhin et al. (2020); Khattab and Zaharia (2020).
@@ 2.3 Text Labeling Methods
An alternative to direct retrieval is to enrich repositories with semantic labels. Unsupervised approaches, such as topic modeling and embedding-based clustering, attempt to discover latent structure without annotation. Classical topic models struggle with short compliance texts Blei et al. (2003), while embedding-based clustering typically assigns each item to a single cluster, limiting representation of overlapping compliance concepts Grootendorst (2022).
Weak supervision generates labels using heuristics or labeling functions Ratner et al. (2017), but requires a predefined label space and substantial upfront design effort. More recently, Large Language Models (LLMs) have been used to generate human-readable labels without fixed taxonomies Tan et al. (2024). However, LLM-based labeling is costly, sensitive to prompt design, and variable across runs Huang et al. (2023). Hybrid approaches reduce some of these limitations but often still rely on repeated LLM inference, constraining scalability in large or evolving repositories.
3 Method
@@ 3.2 Annotation Phase
To form semantically coherent groups without forcing exclusive assignments, we apply Possibilistic C-Means (PCM) Barni et al. (1996); Krishnapuram and Keller (1993). Unlike hard clustering, PCM assigns each question an independent membership score for each cluster, allowing overlap between groups. Let $k$ denote the number of clusters (we set k=|Q|1k=|Q|-1). PCM returns a possibility membership matrix
@@
To convert soft memberships into discrete (possibly overlapping) clusters, we threshold memberships on a per-cluster basis. Specifically, for each cluster $c$, we select a membership threshold $\tau_{c}$ automatically using elbow (knee) detection on the sorted membership scores $\{P_{c,i}\}_{i=1}^{n}$ Satopaa et al. (2011). This avoids dataset-specific manual tuning and adapts to the skewed similarity distributions often observed in compliance question repositories. When no clear elbow is detected, we conservatively fall back to the minimum membership value to avoid discarding potentially relevant associations.
@@
To improve reproducibility and reduce output variance, we use a structured prompt template (Figure 2) Tan et al. (2024); Chen et al. (2025). The template: (i) assigns a domain role (cybersecurity audit classification), (ii) constrains the task to cluster-common labels rather than question-specific details, (iii) enforces comparison across questions to identify shared themes, and (iv) restricts outputs to short noun-phrase labels with no explanations or predefined taxonomies. For each cluster $\mathcal{C}_{i}$, the LLM returns a label set $L_{i}$.
@@ 3.3 Prediction Phase
The prediction phase assigns labels to a new compliance question using the labeled repository produced in the annotation phase, avoiding LLM inference under normal operation. Because the labeled set is typically small and evolves over time, we use instance-based learning with k-Nearest Neighbors (kNN), a non-parametric method that supports incremental updates without retraining Zhou et al. (2003). Given a new question $q^{\prime}$, we compute its embedding $\mathbf{e}^{\prime}=f(q^{\prime})$ and retrieve the $k$ most similar labeled questions in the embedding space using cosine similarity.
To infer labels from the retrieved neighbors, we use a simple voting scheme. Each neighbor contributes its label set, and each label receives one vote per occurrence among the neighbors Cover and Hart (1967). The predicted label set consists of the label(s) with the maximum vote count, allowing ties. To improve robustness, we suppress labels that appear in only one neighbor. If no label receives at least two votes, we treat $q^{\prime}$ as out-of-distribution and defer labeling to the LLM.
@@ 4.1 Experimental Setup
Label quality evaluation assesses whether generated labels accurately capture question intent and remain reusable across semantically related questions, measuring label correctness, label generalization (abstraction beyond individual instances), and consistency (uniform and interpretable naming). Label quality is assessed using an LLM-as-a-Judge Gu et al. (2024) with explicit scoring criteria, enabling scalable comparison across datasets. In addition to the time consumed using different methods and the tokens spent, we also consider the energy consumed. This can help us take into account a multidimensional evaluation of label generation and their quality.
@@ 4.3 Threats to Validity
Finally, the kNN propagation results are influenced by the degree of overlap between the source and target questionnaires. In this paper, we evaluate kNN in a cross-standard transfer setting where the underlying control taxonomies only partially overlap, which can reduce both correctness and label diversity. An important next step is to test kNN under (i) same-distribution conditions, where new questions are sampled from the same questionnaire/standard as the labeled set, and (ii) incremental real-world conditions, where newly authored questions are added over time to an existing, partially overlapping questionnaire. These additional evaluations would clarify how much of the observed kNN degradation is due to cross-standard mismatch versus the intrinsic

### v2

@@ 1 Introduction
As organizations increasingly rely on third-party providers such as cloud platforms and SaaS vendors, Third-Party Risk Assessment (TPRA) has become a core component of cybersecurity risk management and regulatory compliance. TPRA typically relies on structured questionnaires derived from standards such as ISO/IEC 27001 and NIST [8, 15]. Organizations select relevant questions from internal repositories based on a supplier’s scope and risk profile, evaluate the supplier’s responses, and decide whether to proceed with onboarding or request remediation [11, 16].
While significant effort is devoted to analyzing supplier responses, a major operational bottleneck arises earlier in the process: selecting relevant questions. Large organizations maintain extensive repositories of compliance questions accumulated across standards and assessment campaigns, yet these repositories are typically unstructured and curated manually [14]. As a result, question selection is repetitive, time-consuming, and difficult to scale.
Most automated approaches frame question selection as a retrieval problem and rely on text similarity [23], including cosine similarity over dense embeddings [21]. However, such methods operate purely on semantic similarity and lack explicit knowledge of the underlying control domain (e.g., access control, incident response) or assessment scope (e.g., existence verification, enforcement on critical systems). When these aspects are implicit, retrieval often yields overly generic questions that do not align with the assessment scope.
An alternative is to enrich compliance repositories with semantic labels capturing both control domain and scope, enabling label-based retrieval. However, most repositories lack such labels, and there is no established methodology for deriving them automatically. Existing work has explored automated labeling using topic modeling [9] and clustering-based methods [5]. More recently, Large Language Models (LLMs) have shown strong performance as annotators and can generate human-readable labels without predefined taxonomies [24, 22]. Nevertheless, using LLMs as the primary labeling mechanism for large and evolving repositories is costly and sensitive to prompt variability.
@@ 2.1 Third-Party Risk Assessment Systems
TPRA is commonly implemented within third-party risk management (TPRM) workflows that standardize vendor onboarding, questionnaire-based assessments, remediation tracking, and continuous monitoring [11]. Commercial platforms support these workflows through curated question libraries, predefined mappings to compliance frameworks, and reporting pipelines.
@@ 2.2 Similarity-Based Methods
Similarity-based retrieval ranks questions or controls by textual relevance. Traditional lexical methods depend on vocabulary overlap and are sensitive to phrasing variation [19]. Neural retrieval methods embed queries and documents into dense vector spaces, enabling semantic similarity matching [18, 10, 12].
@@ 2.3 Text Labeling Methods
An alternative to direct retrieval is to enrich repositories with semantic labels. Unsupervised approaches, such as topic modeling and embedding-based clustering, attempt to discover latent structure without annotation. Classical topic models struggle with short compliance texts [2], while embedding-based clustering typically assigns each item to a single cluster, limiting representation of overlapping compliance concepts [5].
Weak supervision generates labels using heuristics or labeling functions [17], but requires a predefined label space and substantial upfront design effort. More recently, Large Language Models (LLMs) have been used to generate human-readable labels without fixed taxonomies [22]. However, LLM-based labeling is costly, sensitive to prompt design, and variable across runs [7]. Hybrid approaches reduce some of these limitations but often still rely on repeated LLM inference, constraining scalability in large or evolving repositories.
3 Hybrid Possibilistic Clustering with LLM-Assisted Labeling
@@ 3.2 Annotation Phase
To form semantically coherent groups without forcing exclusive assignments, we apply Possibilistic C-Means (PCM) [1, 13]. Unlike hard clustering, PCM assigns each question an independent membership score for each cluster, allowing overlap between groups. Let $k$ denote the number of clusters (we set k=|Q|1k=|Q|-1). PCM returns a possibility membership matrix
@@
To convert soft memberships into discrete (possibly overlapping) clusters, we threshold memberships on a per-cluster basis. Specifically, for each cluster $c$, we select a membership threshold $\tau_{c}$ automatically using elbow (knee) detection on the sorted membership scores $\{P_{c,i}\}_{i=1}^{n}$ [20]. This avoids dataset-specific manual tuning and adapts to the skewed similarity distributions often observed in compliance question repositories. When no clear elbow is detected, we conservatively fall back to the minimum membership value to avoid discarding potentially relevant associations.
@@
To improve reproducibility and reduce output variance, we use a structured prompt template (Figure 2) [22, 3]. The template: (i) assigns a domain role (cybersecurity audit classification), (ii) constrains the task to cluster-common labels rather than question-specific details, (iii) enforces comparison across questions to identify shared themes, and (iv) restricts outputs to short noun-phrase labels with no explanations or predefined taxonomies. For each cluster $\mathcal{C}_{i}$, the LLM returns a label set $L_{i}$.
@@ 3.3 Prediction Phase
The prediction phase assigns labels to a new compliance question using the labeled repository produced in the annotation phase, avoiding LLM inference under normal operation. Because the labeled set is typically small and evolves over time, we use instance-based learning with k-Nearest Neighbors (kNN), a non-parametric method that supports incremental updates without retraining [25]. Given a new question $q^{\prime}$, we compute its embedding $\mathbf{e}^{\prime}=f(q^{\prime})$ and retrieve the $k$ most similar labeled questions in the embedding space using cosine similarity.
To infer labels from the retrieved neighbors, we use a simple voting scheme. Each neighbor contributes its label set, and each label receives one vote per occurrence among the neighbors [4]. The predicted label set consists of the label(s) with the maximum vote count, allowing ties. To improve robustness, we suppress labels that appear in only one neighbor. If no label receives at least two votes, we treat $q^{\prime}$ as out-of-distribution and defer labeling to the LLM.
@@ 4.1 Experimental Setup
Label quality evaluation assesses whether generated labels accurately capture question intent and remain reusable across semantically related questions, measuring label correctness, label generalization (abstraction beyond individual instances), and consistency (uniform and interpretable naming). Label quality is assessed using an LLM-as-a-Judge [6] with explicit scoring criteria, enabling scalable comparison across datasets. In addition to the time consumed using different methods and the tokens spent, we also consider the energy consumed. This can help us take into account a multidimensional evaluation of label generation and their quality.
@@ 4.3 Threats to Validity
Finally, the kNN propagation results are influenced by the degree of overlap between the source and target questionnaires. In this paper, we evaluate kNN in a cross-standard transfer setting where the underlying control taxonomies only partially overlap, which can reduce both correctness and label diversity. An important next step is to test kNN under (i) same-distribution conditions, where new questions are sampled from the same questionnaire/standard as the labeled set, and (ii) incremental real-world conditions, where newly authored questions are added over time to an existing, partially overlapping questionnaire. These additional evaluations would clarify how much of the observed kNN degradation is due to cross-standard mismatch versus the intrinsic limitations of nearest-neighbor label propagation in sparse semantic spaces.